# Revisiting the Identification of the Conduct Parameter in Homogeneous Goods Markets

### v2

@@ Abstract
We revisit the identification of the conduct parameter in homogeneous goods markets. lauIdentifying1982 argues that the conduct parameter is not identified if and only if the inverse demand function is separable, except for a specific separable function. This result has been regarded as an extension of the result in bresnahanOligopoly1982 to more general settings. However, we show that Lau’s claim is incorrect and provide a new characterization of the non-identification. Our characterization shows that a demand function with demand rotation instruments is the necessary and sufficient condition for the identification of the conduct parameter. Therefore, our result properly generalizes the role of demand rotation instruments in identifying the conduct parameter, as highlighted by bresnahanOligopoly1982, to more general settings.
@@ 1 Introduction
Measuring market power is an important task in several fields of economics, especially in industrial organization. While there are several measures of market power, markup is one of the most popular measures. In the last decade, the rise of markup has been a hot topic in both macroeconomics and industrial organization (see millerIndustrial2025 for the detailed survey). However, when estimating markup, most papers put restrictions on how firms compete, that is, firm conduct, and hence, the estimation result could be biased.111deloeckerRise2020 use accounting data and estimate markup based on the production function approach under the assumption of imperfect competition in an output market and perfect competition in input markets. dopperRising2021 uses Nielsen data and estimates markup under the assumption of Bertrand competition.
In the literature on empirical industrial organization, to measure markup under a flexible competition model, the conduct parameter approach has been used: porterStudy1983, genesoveTesting1998, and okazakiExcess2022 in homogeneous product markets, and millerUnderstanding2017 and cilibertoDoes2014 in differentiated product markets. In this approach, the researcher considers a generalized marginal revenue function with a conduct parameter, and estimating the conduct parameter is equivalent to estimating markup under a flexible competition model. In addition to measuring markup, recently the conduct parameter approach has received attention. weylPassThrough2013 provides a new insight between the conduct parameter and the pass-through and some empirical studies estimate the conduct parameter as an intermediate object to investigate the pass-through (millerPassthrough2017). beringeRobust2025 study the welfare analysis under imperfect competition and use the conduct parameter approach to get a robust prediction of the welfare.
A central question in this approach is the identification of the conduct parameter. As for the identification in homogeneous goods markets, there are two influential papers. First, bresnahanOligopoly1982 considers the identification of the conduct parameter and the marginal cost function in a model with linear demand and linear marginal cost. He shows that the conduct parameter and the parameters of the marginal cost function can be identified when the inverse demand function includes a demand shifter, called a demand rotation instrument. This demand rotation instrument changes the slope and intercept of the inverse demand function without changing the equilibrium point under the true model. However, the equilibrium point changes under the false model, and hence the conduct parameter can be identified by observing the equilibrium outcomes.
Second, lauIdentifying1982 considers a more general setting and investigates the conditions under which the conduct parameter is not identified. Then, he claims that the shape of the inverse demand function plays the central role in the identification by deriving the sufficient and necessary condition for non-identification:
Claim.
Under some assumptions, the conduct parameter and the marginal cost function are not identified if and only if the inverse demand function is separable in demand shifters except for a specific separable function.
The claim implies that separability is the key to the non-identification, but there is a specific separable function that always leads to the identification. The claim also implies the equivalence between non-separable demand functions and the identification. As the inverse demand function with a demand rotation instrument is not separable, this result has been regarded as an extension of bresnahanOligopoly1982’s idea in a more general setting.
The main contribution of our paper is to show that Lau’s claim is incorrect and to provide a new characterization of non-identification. Based on our characterization, we have the following result:
Result.
Under some assumptions, the conduct parameter and the marginal cost function are not identified if and only if inverse demand functions are such that demand shifters can change only the slope or the intercept of the inverse demand function, except for a specific inverse demand function.
In this result, as the demand shifters can change only the slope or the intercept of the inverse demand, the inverse demand function cannot have any demand rotation instrument that requires keeping the equilibrium quantity unchanged by changing the slope and the intercept simultaneously. Therefore, our result restricts the type of the inverse demand function for the non-identification more than the separability requirement in Lau’s claim. The contraposition of our result implies that the inverse demand function with a demand rotation instrument is the necessary and sufficient condition for the identification, and hence our result clearly reinforces the role of demand rotation instruments in the identification of the conduct parameter and indicates that the idea in bresnahanOligopoly1982 is crucial for the identification of the conduct parameter even in more general settings.
Why is Lau’s result incorrect? The idea of Lau’s proof is the following: let $\mathcal{E}=(\theta,MC)$ be the pair of the conduct parameter and the marginal cost function, and let $\mathcal{E}^{*}=(\theta^{*},MC^{*})$ be an alternative pair. The non-identification implies that given demand shifters and cost shifters, both $\mathcal{E}$ and $\mathcal{E}^{*}$ lead to the same equilibrium quantity. Lau shows that the non-identification implies that there is a transformation for the equilibrium conditions between $\mathcal{E}$ and $\mathcal{E}^{*}$, that is, an equilibrium quantity satisfied in $\mathcal{E}$ is also satisfied in $\mathcal{E}^{*}$, which implies that observational equivalence holds. However, we point out that Lau’s transformation does not work properly. When a demand rotation instrument varies so as not to change the equilibrium in $\mathcal{E}$, to keep observational equivalence, his transformation requires that the marginal cost function $MC^{*}$ should depend on the demand shifter, which violates the exclusive demand shifters assumption. Therefore, Lau’s transformation does not work properly.
To resolve the problem in Lau’s transformation, we construct a new transformation of the marginal cost function between $MC$ and $MC^{*}$ based on the definition of the non-identification. Under our transformation, the transformed marginal cost function still could depend on demand shifters. Thus, to make the transformation valid, we restrict the functional form of the inverse demand function such that the transformation is independent of the demand shifter. The restriction leads to a differential equation of the inverse demand function, and the solution of the differential equation leads to the inverse demand function without demand rotation instrument.
Another contribution of our paper is that Lau’s claim assumes that the inverse demand function is already identified. However, we show that the inverse demand function that leads to the non-identification of the conduct parameter and the marginal cost function is not identified. In other words, when the observed data is generated by the inverse demand function that leads to the non-identification of the conduct parameter and the marginal cost function, we cannot identify the inverse demand function, and hence, all model components of the conduct parameter model cannot be identified. In contrast, we show that demand rotation instruments do not help to identify the inverse demand function. Therefore, we conclude that when an inverse demand function with demand rotation instruments is identified, the conduct parameter and the marginal cost function can be identified.
@@
The rest of the paper is organized as follows. Section 2 describes the setting. Section 3 discusses Lau’s result and our main result. Subsection 3.3 and Subsection 3.4 provide a new characterization of non-identification and show the necessary and sufficient condition for non-identification. Section 4 considers whether the inverse demand function that leads to non-identification can be identified. Section 5 discusses some criticisms of the conduct parameter approach. Section 6 concludes. Appendix includes the omitted proofs in the main text and a summary of goldmanNote1964.
@@ 2.1 Conduct Parameter Model
Consider a homogeneous product market with an aggregate inverse demand and an aggregate marginal cost function denoted as $P(Q,\bm{X}^{d})$ and $MC(Q,\bm{X}^{s})$, respectively, where $Q$ is the aggregate product quantity, $\bm{X}^{d}$ and $\bm{X}^{s}$ are the vector of demand shifters and cost shifters, respectively. Note that every vector is represented by a bold letter, and its element or scalar value is represented by a non-bold letter. Let $K_{d}$ and $K_{s}$ be the dimension of $\bm{X}^{d}$ and $\bm{X}^{s}$, respectively.
Given the demand shifter $\bm{X}^{d}$ and the cost shifter $\bm{X}^{s}$, the equilibrium quantity $Q^{e}$ is defined as the solution to the following equation:
@@
where $\varepsilon^{d}$ is the price elasticity of demand. Therefore, the conduct parameter is also regarded as the elasticity-adjusted Lerner index and represents the degree of market power of the firms. Depending on the value of $\theta$, the condition can represent the first-order condition of several models between perfect competition ($\theta=0$) and joint-profit maximization ($\theta=1$). [^2] It can also nest Cournot competition when $\theta=1/N$ under some marginal cost function such as constant marginal cost and linear marginal cost. In general, this holds when the aggregation of each firm’s first-order condition results in the aggregate first-order condition. Therefore, the left-hand side is regarded as a generalized marginal revenue with $\theta$, and hence the condition is a generalized first-order condition with the conduct parameter.
@@
The demand shifter $\bm{X}^{d}$ and the cost shifter $\bm{X}^{s}$ are exclusive; that is, there is no common variable in $\bm{X}^{d}$ and $\bm{X}^{s}$, and $\bm{X}^{d}$ affects the inverse demand function but not the marginal cost function, and $\bm{X}^{s}$ affects the marginal cost function but not the inverse demand function.
We can include some common variables in $\bm{X}^{d}$ and $\bm{X}^{s}$, but in that case, we must have at least one exclusive demand shifter and one exclusive cost shifter. Hereafter, we assume that there is no common variable in $\bm{X}^{d}$ and $\bm{X}^{s}$, for simplicity. The next assumption restricts the effectiveness of the shifters:
@@
for any $Q$, $\bm{X}^{d}$, and $\bm{X}^{s}$.
When this condition is not met, there could be areas where the inverse demand function and the marginal cost function cannot be identified. For example, when a demand shifter does not change the inverse demand function for some interval of the demand shifter, the equilibrium quantity is not affected by the change in the demand shifter. Then, there is no variation to identify the inverse demand function on the interval. Because our identification result for the conduct parameter and the marginal cost function presumes the identification of the inverse demand function. Thus, to guarantee the global identification of the inverse demand function, the assumption is necessary.
Next, we impose the following assumption on the differentiability of the inverse demand function and the marginal cost function:
@@
As we will see, lauIdentifying1982 assumes twice-continuous differentiability of the inverse demand function and the marginal cost function. Hence our result requires a stronger assumption on the inverse demand function. However, we will justify our assumption by showing that Lau’s claim also needs three-times continuous differentiability.
@@
Given an inverse demand function, a conduct parameter, and a marginal cost function, and given $\bm{X}^{d}$ and $\bm{X}^{s}$, the derivative of the equilibrium condition with respect to $Q$ is not zero, that is,
$(1+\theta)\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})+\theta Q\frac{\partial^{2}P}{\partial Q^{2}}(Q,\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q,\bm{X}^{s})$ $\neq 0.$ (4)
@@ 2.2 The Data Generation Process
The identification of the reduced-forms can be done simply due to the variation in the demand shifters and the cost shifters.
2.3 Definitions of Non-identification and Identification
Now, we introduce the definition of the identification problem in this setting. lauIdentifying1982 considers the identification of the conduct parameter and the marginal cost function. While our interest is the identification, we take an indirect approach and specify the conditions under which the model is not identified. The definition of the non-identification is as follows:
@@ 2.3 Definitions of Non-identification and Identification
The conduct parameter and the marginal cost function are said to be non-identified if there are two distinct pairs of conduct parameter and marginal cost function, denoted as $(\theta,MC)$ and $(\theta^{*},MC^{*})$ such that the corresponding reduced-form equilibrium quantity are identical for demand shifter and cost shifter: $h_{q}(\bm{X}^{d},\bm{X}^{s})=h_{q}^{*}(\bm{X}^{d},\bm{X}^{s})$ for any $\bm{X}^{d}$ and $\bm{X}^{s}$. In other words, for each $\bm{X}^{d}$ and $\bm{X}^{s}$, there is an equilibrium quantity $Q^{e}$ that satisfies both
and
$P(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})$ $=MC^{*}(Q^{e},\bm{X}^{s}).$ (7)
Then, by taking the contraposition of Definition 1, we can characterize the identification of the conduct parameter and the marginal cost function:
@@
Note that the non-identification and the identification assume the identification of the inverse demand function. In other words, we consider the identification of the conduct parameter and the marginal cost function given an identified inverse demand function. While it should be discussed whether the inverse demand function can be identified from the data, we will discuss the identification of the inverse demand function later.
@@ 3.1 Lau’s Claim
Based on the setting in Section 2, lauIdentifying1982 shows the equivalence between the separability of the inverse demand function and the non-identification of the conduct parameter and the marginal cost function. The following claim is the quote of Theorem 1 in lauIdentifying1982.
@@
The result says that the shape of the inverse demand function is important for the identification. When the inverse demand function is separable in demand shifters, the conduct parameter is not identified. An inverse demand function satisfying (8) has weak separability defined in goldmanNote1964.  [^3] See Definition B.1 in Appendix B. Note that the separability in goldmanNote1964 is defined only when the dimension of the demand shifter is greater than two. When the demand shifter is a scalar, we cannot apply the definition of separability in goldmanNote1964.
To understand the intuition of Lau’s claim, let us consider an example in bresnahanOligopoly1982. In his paper, he considers a market with a linear inverse demand and a linear marginal cost. To break the separability, he considers a linear inverse demand function such that
@@
where $X^{d}_{1}$ is called a demand rotation instrument because it can change the slope and the intercept of the inverse demand function without changing the equilibrium quantity. While Bresnahan shows the identification under the linear inverse demand function, Lau’s claim extends Bresnahan’s result to more general settings.
The claim also emphasizes that the dimension of the demand shifter is important for the identification. When the demand shifter is a scalar, that is, when the dimension of the demand shifter is one, the inverse demand function (8) nests any inverse demand with a scalar demand shifter because under a scalar demand shifter $X^{d}$, the change of the demand shifter through $r(X^{d})$ is regarded as the same as the change in $X^{d}$. Thus, the conduct parameter can be identified only when the inverse demand function takes the form (9) when the demand shifter is a scalar.
While Lau’s claim has been regarded as the extension of bresnahanOligopoly1982 to more general settings, an important observation is that introducing demand rotation instruments breaks the separability, but the separability does not exclude demand rotation instruments. For example, consider an inverse demand function given as
@@
where the range of $r$ is greater than one.
First, based on the definition of weak separability in goldmanNote1964, the function is verified to be separable because for any $i$ and $j$:
$\frac{\partial}{\partial Q}\left(\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})}{\frac{\partial P}{\partial X^{d}_{j}}(Q,\bm{X}^{d})}\right)=\frac{\partial}{\partial Q}\left(\frac{r_{i}(\bm{X}^{d})(1-Q)}{r_{j}(\bm{X}^{d})(1-Q)}\right)=\frac{\partial}{\partial Q}\frac{r_{i}(\bm{X}^{d})}{r_{j}(\bm{X}^{d})}=0,$ (12)
@@
Second, we can verify that the demand shifters can work as demand rotation instruments. Given the inverse demand function, the marginal revenue under $\theta$ is written as
@@
When the quantity is $Q^{\prime}=\frac{1}{1+\theta}$, the marginal revenue is not affected by the change in the demand shifter $X^{d}$ because it is equal to one. Suppose that there is a marginal cost function $MC$ where $Q^{\prime}$ holds as an equilibrium. Then, as the marginal revenue is a constant in the demand shifter, any change in the demand shifter does not affect the equilibrium quantity. In contrast, suppose that under another conduct parameter $\theta^{*}$ and another marginal cost function $MC^{*}$, $Q^{\prime}$ holds as an equilibrium, that is, $Q^{\prime}$ satisfies
$r(\bm{X}^{d})(1-(1+\theta^{*})Q^{\prime})+Q^{\prime}(1+\theta^{*})=MC^{*}(Q^{\prime},\bm{X}^{s}).$ (14)
Here, the marginal revenue under $\theta^{*}$ is a function of the demand shifter at $Q^{\prime}$. Thus, the change in $\bm{X}^{d}$ leads to a different equilibrium quantity from $Q^{\prime}$, which implies the violation of the observational equivalence. Therefore, the demand shifter $\bm{X}^{d}$ can work as the demand rotation instrument at $Q^{\prime}$ because the demand shifter changes the slope and the intercept of the inverse demand function simultaneously without changing the equilibrium quantity under the true conduct but the change leads to a different equilibrium quantity under false conduct.
While the example admits a separable inverse demand function with demand rotation instrument, unfortunately, Lau does not clearly explain how this situation is treated in his claim, and hence, it is not clear why not demand rotation instruments but the separability of the inverse demand function is the key to identification.

### v3

# Yuri Matsumura    Suguru Otani

@@ Abstract
We revisit the identification of the conduct parameter in homogeneous goods markets. Lau (1982) shows that the conduct parameter is not identified if and only if the inverse demand function is separable, except for a specific separable function. This result has been regarded as an extension of the result in Bresnahan (1982) to more general settings. However, we show that Lau’s claim is incorrect and provide a new characterization of the non-identification. Our characterization shows that the presence of demand rotation instruments in the demand function is the necessary and sufficient condition for identifying the conduct parameter. Therefore, our result properly generalizes the role of demand rotation instruments in identifying the conduct parameter, as highlighted by Bresnahan (1982), to more general settings.
@@ 1 Introduction
Empirical research in industrial organization examines a wide range of questions, including the measurement of market power and concentration, the evaluation of welfare, the assessment of cartels, and the prediction of merger impacts. In virtually all of these applications, empirical conclusions depend critically on assumptions about how firms compete, that is, on the specification of firm conduct. Researchers often impose a particular competitive model, such as price-taking, Cournot, or Bertrand, because conduct is rarely identified without additional information, such as detailed cost data or industry- or institution-specific information. Nevertheless, misspecifying conduct can distort the measurement or lead to misleading conclusions.
To address this problem, the conduct parameter approach has been used in the literature on industrial organization. This approach embeds a scalar measure of firm conduct into the marginal revenue function and allows researchers to estimate conduct directly from data, providing a flexible model to investigate several questions. The applications vary across industries and questions, including homogeneous-product markets (Porter 1983, Genesove and Mullin 1998, Okazaki, Onishi, and Wakamori 2022), differentiated-product markets (Miller and Weinberg 2017, Ciliberto and Williams 2014, Sullivan 2020), pass-through analysis (Weyl and Fabinger 2013, Miller, Osborne, and Sheu 2017), welfare evaluation under imperfect competition (Beringe and Whitmeyer 2025, Nocke and Schutz 2025), and merger evaluation (Aryal, Chattopadhyaya, and Ciliberto 2025).
Despite its broad use, the basic identification result on the conduct parameter in a homogeneous goods market remains rooted in two classic contributions. First, Bresnahan (1982) shows that, with a linear demand and a linear marginal cost, a demand rotation instrument can identify the conduct parameter by moving the slope and intercept of inverse demand simultaneously because it can keep the equilibrium the same under the true conduct parameter while changing the equilibrium under the false conduct parameter. Second, Lau (1982) considers an environment with a more general inverse demand function and a marginal cost function and shows that conduct is not identified if and only if the inverse demand function is separable in demand shifters. Because demand rotation instruments break the separability of the inverse demand function, Lau’s condition has been interpreted as a general characterization of identification and regarded as an extension of Bresnahan’s idea in general settings.
The main contribution of this paper is to show that Lau’s characterization is incomplete and to provide a new condition that fully characterizes if and only if conduct and marginal cost are not identified. Our main result shows that the non-identification of the conduct parameter and the marginal cost function arises if and only if demand shifters can change only the slope or only the intercept of the inverse demand function. In this case, the inverse demand function cannot have any demand rotation instrument, and hence conduct is identified precisely when the inverse demand function includes demand rotation instruments. This result is both simpler and more restrictive than separability in Lau’s claim and clarifies the precise role of demand rotation instruments in identifying the conduct parameter in general environments.
The intuition is the following. Given a conduct parameter and a marginal cost function, the non-identification implies that we can transform the marginal cost function into another marginal cost function that, together with another conduct parameter, produces the same equilibrium quantity. However, this transformation is usually invalid because keeping the equilibrium points the same in both models requires the transformed marginal cost function to depend on demand shifters, which violates the exclusive demand shifters assumption. This dependence disappears only when demand shifters can change just the slope or just the intercept of the inverse demand function, in which case we can construct different pairs of conduct parameters and marginal cost functions that yield the same equilibrium for any value of the shifters without violating the exclusion restriction. Thus, when demand rotation instruments are present in the inverse demand function, the transformation remains invalid, preventing observational equivalence between alternative conduct and marginal-cost pairs and ensuring identification. The issue that makes Lau’s proof incomplete is that while he also derives a transformation between marginal cost functions, he fails to eliminate the dependence of the transformed marginal cost function on demand shifters.
@@
Taken together, our results refine the theoretical foundations of conduct parameter estimation and clarify the exact role that demand shifters must play to allow practitioners to implement empirical analysis with a flexible competition model. The rest of the paper is organized as follows. Section 2 describes the setting. Section 3 discusses Lau’s result and our main result. Subsections 3.3 and 3.4 provide a new characterization of non-identification and show the necessary and sufficient condition for non-identification. Section 4 discusses some criticisms of the conduct parameter approach. Section 5 concludes. The Appendix includes the omitted proofs in the main text and a summary of Goldman and Uzawa (1964).
@@ 2.1 Conduct Parameter Model
Consider a homogeneous product market with an aggregate inverse demand and an aggregate marginal cost function denoted by $P(Q,\bm{X}^{d})$ and $MC(Q,\bm{X}^{s})$, respectively, where $Q$ is the aggregate product quantity, $\bm{X}^{d}$ and $\bm{X}^{s}$ are the vectors of demand shifters and cost shifters, respectively. Note that every vector is represented by a bold letter, and its element or scalar value is represented by a non-bold letter. Let $K_{d}$ and $K_{s}$ be the dimension of $\bm{X}^{d}$ and $\bm{X}^{s}$, respectively.
Given the demand shifter $\bm{X}^{d}$ and the cost shifter $\bm{X}^{s}$, the equilibrium quantity $Q^{e}$ solves the following equation:
@@
where $\varepsilon^{d}$ is the price elasticity of demand. Therefore, the conduct parameter is also regarded as the elasticity-adjusted Lerner index and represents the degree of market power of the firms. Depending on the value of $\theta$, the condition can represent the first-order condition of several models between perfect competition ($\theta=0$) and joint-profit maximization ($\theta=1$). [^1] It can also nest Cournot competition when $\theta=1/N$ under some marginal cost function such as constant marginal cost and linear marginal cost. In general, this holds when the aggregation of each firm’s first-order condition results in the aggregate first-order condition. Therefore, the left-hand side of (1) is regarded as a generalized marginal revenue with $\theta$, and hence the condition is a generalized first-order condition with the conduct parameter.
@@
The set of all exogenous variables affecting the market equilibrium can be partitioned into (1) exclusive demand shifters, $\bm{X}^{d}$, that affect the inverse demand function but not the marginal cost function, (2) exclusive cost shifters, $\bm{X}^{s}$, that affect the marginal cost function but not the inverse demand function, and (3) common shifters, $\bm{Z}$, that affect both the inverse demand function and the marginal cost function.
The theoretical analysis of identification often simplifies the setting by assuming that the demand shifters and cost shifters are mutually exclusive, as per Assumption 1. We formalize our approach by conditioning the analysis on an arbitrary realization $\bm{z}$ of these common shifters. This means all subsequent functions are understood to be conditional functions, $P(Q,\bm{X}^{d})\equiv\tilde{P}(Q,\bm{X}^{d},\bm{Z}=\bm{z})$ and $MC(Q,\bm{X}^{s})\equiv\tilde{MC}(Q,\bm{X}^{s},\bm{Z}=\bm{z})$. By treating $\bm{Z}$ as fixed constants in this conditional analysis, we maintain the notational simplicity of focusing only on the exclusive shifters while ensuring that the derived non-identification characterization holds robustly across all values of the common variables in empirical settings.
The next assumption restricts the effectiveness of the shifters:
@@
for any $Q>0$, $\bm{X}^{d}$, and $\bm{X}^{s}$.
When this condition is not met, there could be areas where the inverse demand function and the marginal cost function cannot be identified. For example, when a demand shifter does not change the inverse demand function for some interval of the demand shifter, the equilibrium quantity is not affected by the change in the demand shifter. Then, there is no variation to identify the inverse demand function on the interval. Because our identification result for the conduct parameter and the marginal cost function presumes the identification of the inverse demand function, to guarantee the identification of the inverse demand function, the assumption is necessary.
Next, we impose an assumption on the differentiability of the inverse demand function and the marginal cost function:
@@
As we will see, Lau (1982) only assumes twice-continuous differentiability of the inverse demand function and the marginal cost function. Hence, our result requires a stronger assumption on the inverse demand function. However, we will justify our assumption by showing that Lau’s claim also needs three-times continuous differentiability later.
@@
Given an inverse demand function, a conduct parameter, and a marginal cost function, and given $\bm{X}^{d}$ and $\bm{X}^{s}$, the derivative of the equilibrium condition with respect to $Q$ at the equilibrium quantity $Q^{e}$ is not zero, that is,
$(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s})$ $\neq 0.$ (4)
@@ 2.2 The Data Generation Process
The identification of the reduced forms follows directly from variation in the demand and cost shifters.
Remark 1.
While Bresnahan (1982) considers a model with error terms, by following Lau (1982), we do not consider any error term in the demand and the supply side, that is, our model does not have any unobserved characteristics to the econometrician. This is a restrictive assumption because there could be real data where $\bm{X}^{d}$ and $\bm{X}^{s}$ are the same in two time periods, but the equilibrium outcomes are different in the two periods. As the scope of this paper is to fully characterize the identification condition along with Lau’s setting,. In other words, our identification analysis is independent of the presence of error terms and concerns whether the structural mapping from primitives $(\theta,MC)$ to the reduced-form equilibrium outcomes $(h_{q},h_{p})$ is injective, that is, whether $(\theta,MC)$ can be uniquely recovered given $(h_{q},h_{p})$. We do not include unobserved characteristics and leave it for future research.
2.3 Definitions
Now, we introduce the definition of the identification problem in this setting. While our interest is the identification of the conduct parameter and the marginal cost function, by following Lau (1982), we take an indirect approach and specify the conditions under which the model is not identified. Our definition of non-identification is based on when two different models lead to observational equivalence:
@@ 2.3 Definitions
The conduct parameter and the marginal cost function are said to be non-identified if there are two distinct pairs of conduct parameters and marginal cost functions, denoted by $(\theta,MC)$ and $(\theta^{*},MC^{*})$, such that the corresponding reduced-form equilibrium quantity that satisfies the equilibrium condition (1) and the corresponding reduced-form equilibrium price are identical for every demand- and cost-shifter value:
$Q^{e}=h_{q}(\bm{X}^{d},\bm{X}^{s})=h_{q}^{*}(\bm{X}^{d},\bm{X}^{s})\text{ and }P^{e}=h_{p}(\bm{X}^{d},\bm{X}^{s})=h_{p}^{*}(\bm{X}^{d},\bm{X}^{s})$
for any $\bm{X}^{d}$ and $\bm{X}^{s}$.
By taking the contraposition of Definition 1, we can characterize the identification of the conduct parameter and the marginal cost function:
@@
Note that the non-identification and the identification assume that the inverse demand function is already identified. In other words, we study the identification of the conduct parameter and the marginal cost function given an identified inverse demand function.
@@ 3.1 Lau’s Claim
Lau (1982) investigates the identification of the conduct parameter and the marginal cost function and obtains the following claim, which is the quote of Theorem 1 in Lau (1982):
@@
The result states that the separability of the inverse demand function is crucial for the identification, but there is a type of separable inverse demand function that can lead to the identification. An inverse demand function satisfying (6) has weak separability defined in Goldman and Uzawa (1964). [^2] See Definition B.1 in Appendix B. Note that the separability in Goldman and Uzawa (1964) is defined only when the dimension of the demand shifter is greater than two. When the demand shifter is a scalar, we cannot apply the definition of separability in Goldman and Uzawa (1964). The claim also emphasizes that the dimension of the demand shifter is important for the identification. When the demand shifter is a scalar, (6) nests any inverse demand function because we can set $\tilde{P}(Q,\bm{X}^{d})=P(Q,r(\bm{X}^{d}))$. Thus, with a scalar demand shifter, the conduct parameter can be identified only when the inverse demand function takes the form (7).
To understand the intuition of Lau’s claim, let us consider an example in Bresnahan (1982) without an error term. Bresnahan considers a market with a linear inverse demand and a linear marginal cost, where the linear inverse demand function is given as
@@
where $X^{d}_{1}$ is a demand rotation instrument because it can change the slope and the intercept of the inverse demand function without changing the equilibrium quantity. It is easy to verify that the demand rotation instrument breaks the separability of the inverse demand function. Additionally, the dimension of the demand shifter is greater than one, and hence Lau’s claim implies that even a nonlinear marginal cost function, (8), leads to the identification of the conduct parameter.
While Lau’s claim has been regarded as an extension of Bresnahan (1982) to more general settings, an important observation is that including demand-rotation instruments is inconsistent with imposing separability on the inverse demand function, but the separability assumption itself does not preclude the existence of demand-rotation shifters in an inverse demand function.
For example, consider an inverse demand function given as
@@
where the range of $r$ is greater than one, that is, $r(\bm{X}^{d})\geq 1$ for any $\bm{X}^{d}$. Assume that the dimension of $\bm{X}^{d}$ is greater than two. Note that in this inverse demand function, the demand shifter can change the slope and the intercept of the inverse demand function.
First, the function is verified to be separable because for any $i$ and $j$, we have
$\frac{\partial}{\partial Q}\left(\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})}{\frac{\partial P}{\partial X^{d}_{j}}(Q,\bm{X}^{d})}\right)=\frac{\partial}{\partial Q}\left(\frac{r_{i}(\bm{X}^{d})(1-Q)}{r_{j}(\bm{X}^{d})(1-Q)}\right)=\frac{\partial}{\partial Q}\left(\frac{r_{i}(\bm{X}^{d})}{r_{j}(\bm{X}^{d})}\right)=0,$ (10)
@@
Second, we can verify that the demand shifters can work as demand rotation instruments to break observational equivalence. Given the inverse demand function, the marginal revenue under $\theta$ is written as
@@
When the quantity is $Q^{\prime}=\frac{1}{1+\theta}$, the marginal revenue is not affected by the change in the demand shifter $\bm{X}^{d}$ because it is equal to one. Suppose that there is a marginal cost function $MC$ where $Q^{\prime}$ holds as an equilibrium for some $\tilde{\bm{X}^{s}}$. Then, as the marginal revenue is a constant in the demand shifter, any change in the demand shifter does not affect the equilibrium quantity. In contrast, suppose that under another conduct parameter $\theta^{*}$ and another marginal cost function $MC^{*}$, $Q^{\prime}$ holds as an equilibrium for some $\tilde{\bm{X}^{d}}$ and $\tilde{\bm{X}^{s}}$ , that is, $Q^{\prime}$ satisfies
[…]
Here, the marginal revenue under $\theta^{*}$ is a function of the demand shifter at $Q^{\prime}$. Thus, the change in $\bm{X}^{d}$ leads to a different equilibrium quantity from $Q^{\prime}$, which implies the violation of the observational equivalence. Therefore, the demand shifter $\bm{X}^{d}$ can work as the demand rotation instrument at $Q^{\prime}$ because it changes the slope and the intercept of the inverse demand function simultaneously without changing the equilibrium quantity under the true conduct, whereas the change leads to a different equilibrium quantity under any false conduct.
Furthermore, when we treat $r(\bm{X}^{d})$ as a scalar demand shifter, the second argument implies that even a scalar demand shifter can break observational equivalence, although the inverse demand function does not satisfy (7). The example admits a separable inverse demand function with a demand rotation instrument, but Lau does not clearly explain how this situation fits his claim, and hence it is not clear why separability, rather than the availability of demand rotation instruments, is the key to identification.